OpenStudy (anonymous):

-5x + 4(2x + 2) = -1 I need some help!

OpenStudy (anonymous):

So far I have and correct me if I am wrong. -5x + 4(2x + 2) = -1 -5x + 2x + 8 = -1 (-5x + 2x) = (-1 + 8) -3x = 7 ????

OpenStudy (shamim):

No

OpenStudy (shamim):

-5x+8x+8=-1

OpenStudy (shamim):

3x=-9

OpenStudy (shamim):

Do the rest

OpenStudy (shamim):

I think u got it

OpenStudy (anonymous):

Thank You!